### v1

# LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues

Abstract
Virtual assistants are poised to take a dramatic leap forward in terms of their dialogue capabilities, spurred by recent advances in transformer-based Large Language Models (LLMs). Yet a major bottleneck to achieving genuinely transformative task-oriented dialogue capabilities remains the scarcity of high quality and linguistically sophisticated data. Existing datasets, while impressive in scale, have limited domain coverage and contain few genuinely challenging conversational phenomena; those which are present are typically unlabelled, making it difficult to assess the strengths and weaknesses of models without time-consuming and costly human evaluation. Moreover, creating high quality dialogue data has until now required considerable human input, limiting both the scale of these datasets and the ability to rapidly bootstrap data for a new target domain. We aim to overcome these issues with LUCID, a modularised and highly automated LLM-driven data generation system that produces realistic, diverse and challenging dialogues. We use LUCID to generate a seed dataset of 4,277 multi-domain, multi-intent conversations across 100 intents to demonstrate its capabilities. The generated conversations include a wide range of challenging phenomena and diverse user behaviour, identifiable via a set of turn-level tags. Finally, we provide separate test sets for seen and unseen intents, allowing for convenient out-of-distribution evaluation. We release both the data generation code and the dataset itself [^1] Please see http://github.com/apple/ml-lucid-datagen for the data and project code.
LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues
[…]
1 Introduction
Figure 1: An extract of a LUCID conversation containing a challenging phenomenon. In this case, the second user response is most likely to be from an overheard conversation rather than providing the desired slot value.
As AI virtual assistants become more sophisticated, there will be an increasing need for dialogue datasets with more challenging conversational phenomena that can be used for model fine-tuning and evaluation. Existing datasets include multi-turn, multi-intent and multi-domain conversations Rastogi et al. (2020); Budzianowski et al. (2018), in addition to multi-lingual datasets Goel et al. (2023); FitzGerald et al. (2023); Hung et al. (2022); Li et al. (2021). However, in each case, the number of intents covered is relatively small. Moreover, the conversational phenomena included in these datasets are often limited in scope. The PRESTO dataset (Goel et al., 2023) does explicitly label specific conversational phenomena, but to the best of our knowledge it is unique in this respect. Additionally, current machine-to-machine data collection methods still involve varying degrees of human involvement, with humans paraphrasing machine generated templates into natural language, and/or manually crafting plausible sequences of intents as dialogue outlines (Shah et al., 2018; Rastogi et al., 2020). Even this reduced level of human involvement in the data generation process, relative to even more onerous human-to-human or human-to-machine approaches, is costly and time consuming, and difficult to scale to new target domains.
To overcome these issues, we introduce LUCID, LLM-generated Utterances for Complex and Interesting Dialogues. LUCID is composed of a pipeline of modularised LLM calls that create realistic, accurate and complex data, allowing the data generation process to scale to more intents, slots and challenging conversational phenomena. In particular, based on a simple description of an intent (for example ‘send an email’), LUCID automatically generates a schema for the intent, including a list of required and optional slot names, types and values. A mock back-end [^2] The mock back-end converts intents into Python classes, which are then instantiated as objects whose interface is exposed to the system agent is then created for the intent that interacts with LLM-driven system and user agents. The system agent produces system commands, which are then executed by the mock back-end.
To introduce diversity and complexity into the data generation process, conversation rules are provided to the user agent to guide it towards specific conversational phenomena. Adding new conversational phenomena to the system is a simple matter of adding a set of imperatives written in natural language to a python file, for example "When Lucid asks about {arg1}, don’t answer, but instead say something to another person who is in the same room" (see Figure 1).
Ensuring data quality is a central challenge for a machine-to-machine generation process. LUCID addresses this issue by breaking down the generation process into a pipeline of multiple, simpler LLMs calls, thereby compartmentalising the data generation task into manageable steps that an LLM can consistently perform accurately. In addition, we use multiple LLM-based validators which discard conversations that might contain an issue. Our if in doubt, discard philosophy ensures a high quality standard for the data being created. Finally, we use sampling throughout our generation process to ensure variety in the length and content of each conversation. To showcase the quality, realism and sophistication of the data generated by LUCID, we release a seed dataset of 92,699 turns.
To summarise our contributions: 1) We introduce a highly automated methodology for producing high quality LLM-generated conversations. 2) We provide a challenging dataset of 4,277 dialogues across 100 intents, including more labelled conversational phenomena than previous dialogue datasets. 3) We provide two tests sets, a test set for seen intents, and an additional test set for unseen intents, allowing for convenient out-of-distribution evaluation. 4) We release the data generation code to enable large scale data generation across different intents and domains, with the option of adding additional, complex conversational flows.
2 Related Work
2.1 Task Oriented Dialogue Datasets
The most popular approach for creating dialogue datasets involves human-to-human interactions, with user annotators interacting with Wizard of Oz (WoZ) annotators Budzianowski et al. (2018); El Asri et al. (2017); Zhu et al. (2020); Eric et al. (2017); Wen et al. (2017). While using human annotators can create diverse, large scale datasets, this is done at a considerable cost, with expert annotators required for accurate dialogue annotations. User annotators follow a generated conversation plan (Budzianowski et al., 2018; El Asri et al., 2017; Zhu et al., 2020), guiding their interactions with the WoZ agent. These instructions can include how the user should navigate specific scenarios, for example when no results match their constraints (El Asri et al., 2017). We find that even in a purely machine-to-machine setup, generating conversation plans for each dialogue remains an effective way to ensure conversational variety. LUCID extends these plans to also specify how and when any complex conversational phenomena should be included in the dialogue.
2.2 Automated Data Collection Methods
To reduce the workload of annotators, dataset collection is becoming increasingly automated. A popular approach is to generate conversation outlines, which are then converted into natural language by annotators Shah et al. (2018); Rastogi et al. (2020); Lin et al. (2021) or using natural language templates (Bordes et al., 2017). As these conversation outlines are simulated based on hard-coded rules, this can limit the diversity of the user behaviour.
Conversations can also be created based on user-machine interactions, for example with users telephoning computerized agents Henderson et al. (2014a, b). Recently, Goel et al. (2023) collected conversations between annotators and a virtual assistant simulator to generate natural language conversation histories, which provide context to semantically annotated user turns.
Human involvement in automated data generation includes ensuring the quality of the dataset, paraphrasing user and agent responses (Shah et al., 2018; Rastogi et al., 2020), providing semantic annotations (Goel et al., 2023; Budzianowski et al., 2018), outlining the sequences of user intents (Rastogi et al., 2020), and identifying out of scope or incoherent examples (Goel et al., 2023). We show that, with recent advances in language modelling OpenAI (2023), by reducing the data generation task into manageable steps, and using our if in doubt, discard validation methodology, it is now possible to achieve the same quality in an almost entirely machine-to-machine generation process.
2.3 Data Generation with LLMs
Wu et al. (2023) recently introduce a framework allowing the interaction of multiple, different LLMs, based on the idea that LLMs can solve highly challenging tasks if these tasks are broken into smaller steps. While Wu et al. (2023) are successful in generating dialogues for a group chat scenario, this does not require the intent and slot labelling needed for task-oriented dialogue. For other NLP tasks, to avoid labelling errors or poor quality data, generating data with LLMs can involve human annotators reviewing the generated utterances Liu et al. (2022); Wiegreffe et al. (2022), or using the generated data as unlabelled data to be used with knowledge distillation Meng et al. (2022); He et al. (2022); Stacey and Rei (2023). Labeled data generation has been successful for other tasks without human input Honovich et al. (2023); Ye et al. (2022); Schick and Schütze (2021); West et al. (2022); Wu et al. (2022), however noise may be an issue for a large proportion of the data Honovich et al. (2023); Schick and Schütze (2021).
3 Method
LUCID decomposes the data generation process into 14 individual LLM calls [^3] At this time we are unable to provide the details of the LLMs we used at each stage. If we receive this approval, we will provide an updated version of the paper., described here as stages, creating manageable steps that LLMs can perform accurately. Alongside our validation process, including our if in doubt, discard approach, reducing the complexity of each LLM call helps to ensure the quality of our generated data. The data generation process consists of four main components (see Figure 2): the generation of intents (stages 1-2), a conversational planner (stages 3-8), turn-by-turn generation of conversations (stages 9-12), and our validation process (stages 13 and 14). The turn-by-turn data generation involves a User LLM agent interacting with a System LLM agent, which in turn interacts with a mock back-end created for each of our generated intents.
Figure 2: The stages in the LUCID data generation, generating intents (stages 1-2), planning conversations (stages 3-8), generating the conversations (stages 9-12) and validating the system predictions (stages 13-14).
3.1 Intent Generation
Schema for each intent are generated by an LLM, using a short human-authored natural language description of the intent (stage 1). In total, we provided LUCID with 54 transactional intent descriptions across 13 domains (see Appendix D). Using these descriptions, LUCID calls an LLM to generate the intent and slot names, as well as the data type of each slot and whether it is mandatory or optional. Each transactional intent has a corresponding query intent [^4] There are some cases when transactional intents have the same query intent - see Appendix D for details, with the purpose of retrieving entities created by that intent.
The next stage (stage 2), involves generating plausible values for each slot. We use these slot values as a starting point for our conversation planner, helping to encourage varied conversations.
3.2 Conversation Planner
The conversation planner provides instructions that guide a user LLM agent down certain types of conversational flows. The planner specifies: 1) the sequence of intents, 2) the slot values for each intent, and 3) any complex conversational phenomena that must be included, specifying when and how these phenomena should be incorporated. This creates a plan that the user simulator must adhere to, reducing the complexity of the data generation task, while also ensuring variety in the generated conversations. This plan is communicated to the User LLM at each turn through a series of conversation rules, specifying the next intent (and its corresponding slot values) that should be requested, in addition to specifying the next complex conversational phenomena. The planner creates the conversation plan at the start of each conversation, involving a pipeline of different stages (stages 3-8).
3.2.1 Generating Conversation Structures
Firstly, the planner decides the sequence of intents that will be included in a conversation. Depending on the primary intent used to start the conversation, the planner then decides which intents are likely to follow this intent, with the aim of creating both varied and realistic conversations.
To achieve this, the planner is provided with a sample of intents across all domains. An LLM then generates a plausible sequence of intents from this sample (stage 3). To ensure variety in the generated conversations, the planner makes further use of sampling, choosing how many intents should be provided in the conversation, which optional slots should be discussed, the conversational phenomena (both happy and unhappy paths) that should be included, and which slots and intents any conversational phenomena should be applied to.
Figure 3: Examples for eight of the nine challenging conversational phenomena included in the LUCID dataset. We also included ‘cancellation’ examples which are similar to ‘delay confirmation’, resulting in the system not confirming a given intent.
3.2.2 Generating Slot Values
The slot values chosen by the planner substantially impact the conversations, and as a result, we have multiple, separate stages for generating the slot values (stages 4, 5, 6 & 7). Choosing slot values that are not plausible or coherent leads to unrealistic conversations. Moreover, to ensure variety in our generated conversations, we require variety within the slot values decided by the planner.
This process involves updating the slot values to make sure these are realistic and coherent (stage 4), generating a reason why the user wants to perform any subsequent intents (stage 5), and generating slot values for the subsequent intents based on this justification (stage 6). Finally, an LLM updates the slot values across every intent in the conversation to ensure they are consistent and realistic when considered collectively (stage 7). See Appendix E for further details.
We use an LLM to generate realistic entities to be returned after any queries. These entities will be provided to the Response LLM to communicate back to the user (stage 8).
3.3 Generating Conversations
Conversations are generated turn-by-turn with a User LLM interacting with a System LLM, which in turn interacts (via Pythonic function calls and variable assignments) with a mock back-end for each intent, and a Response LLM communicating natural language responses back to the user. The user behaviour is governed by the conversational rules created by our conversation planner, shaping the outcome of the conversations.
Conversations start with an utterance from the User LLM (stage 9), which is then interpreted and labelled by the System LLM (stage 10). Initially, no string slot values are predicted. These values are predicted in an additional stage (stage 11), where an LLM is instructed, where possible, to choose the string values from spans of the user utterance, providing fewer opportunities for the LLM to hallucinate. The predicted semantic labels then interact with a mock back-end for each intent, assessing whether any additional slots are required or whether the user should instead be asked to provide confirmation. The mock back-end provides these instructions to a Response LLM (stage 12), which generates a natural language response to the user.
3.4 LLM-based Validation
We implement an LLM-based validation process to ensure reliable and consistent labelling in the conversations, using our principle of if in doubt, discard. First, based on the observation that the system LLM is more uncertain about incorrect predictions, we repeat the system predictions twice (using temperature values less than 1), and abort the conversation if the three predictions are not identical (stage 13). Additional validation is then performed by another LLM (stage 14) which also labels the user requests, except this time with access to the conversation rules that the user is following. Predictions must exactly match the original System LLM outputs, otherwise the conversation is discarded.
Further validation is provided for our conversational phenomena, described in Section 3.5, in addition to a range of post-processing rules introduced after a qualitative analysis (described in Section 4.3). In total, 56% of the conversations generated passed our validation checks (see Appendix H for further information).
To avoid wasting valuable conversational data, we salvage the prefix of an aborted conversation up to the point where the validation error was identified [^5] To avoid overly short conversations, we do this only if at least one intent has been performed already or at least 10 turns have occurred. In these cases, we truncate the conversation, sampling LLM generated natural language responses that justify interrupting the conversation.
3.5 Introducing Additional Conversational Phenomena
To make interesting, diverse and challenging conversations, we introduce a wide range of conversational phenomena which are labelled automatically at the turn level (see Figure 3). These phenomena include sarcastic or irrelevant replies, or cases where the user is overheard in another conversation. LUCID also contains examples where a user corrects themselves, either within a turn (in-turn correction) or in a later turn (correction). Alternatively, a user may cancel an intent (cancellation) or delay confirming the intent until a future turn (delay confirmation). Our conversational phenomena also consider how the user responds to prompts from the virtual assistant. If the virtual assistant requests a value for one slot, the user can respond about a different slot (respond different slot). See Table 3 for the full distribution of these phenomena in our data.
The conversation rules are sampled by the planner for each conversation, before the planner assigns these phenomena to specific intents (and slots) within the conversation. In addition to the unhappy path conversational phenomena discussed, we also introduce different happy paths via conversation rules to introduce variety within conversations. These rules state whether or not the user should provide some, all or no slot values when expressing the intent for the first time.
We introduce additional validations to ensure these challenging turns are correctly predicted by our LLMs. When the conversation rules instruct a user to introduce a specific phenomenon for a certain slot value, the user is instructed to also provide a signal (in the form of a special token) to show that this unhappy path is being performed. We then use this signal for validation purposes, ensuring that the following system command matches the expected response for this phenomenon. However, we do not provide these special tokens to the system which interprets and labels the user request; these would not be available to a real virtual assistant and we find that including them during data generation can result in unrealistic target labels (e.g. if a user’s ‘irrelevant’ response accidentally constitutes a plausible slot value).
Figure 4: An example labeled conversation (a simplified example from the dataset). Each dialogue contains user turns, system, signal and response turns.
3.6 Annotation Scheme and our Mock Back-end
We apply a concise labelling system to track the states for each intent in a conversation. This labelling system follows a Pythonic syntax, with function calls used to initialise intents and entities when these are first mentioned, and attribute assignments used for any subsequent slot filling operations (see Figure 4). This schema-based, concise form of semantic labelling is highly convenient, and avoids the need for state tracking for each individual turn. The Pythonic variables also provide a natural way to encode anaphoric phenomena such as pronominal coreference, as well as effectively leveraging the LLM’s pretraining data, in which Python is likely very well represented.
Our labeling involves four types of turns in a conversation: 1) User turns; 2) System turns, labelling the intentions of a user (including slot values and the intents specified); 3) Signal turns, returned after our mock back-end processes the system command, recommending the next action; and 4) Response turns, which are natural language responses back to the user.
System turns always follow user turns. In most cases, the first system turn is followed by a signal turn, except when the system decides to immediately call a response with ‘say()’, for example if the user response is irrelevant. Signal turns are then followed by a system turn, which triggers the natural language response turn. The system turns therefore decide when to pass information to the mock back-end, and when to trigger the natural language response. We use the system turns as the targets in this dataset.
LUCID automatically creates the mock back-end for each intent using the schema generated in steps 1 and 2. This involves generating a Python class to represent the intent in question, which is then instantiated as an object and interacts with the system commands to indicate when mandatory slots have not yet been provided, or when confirmation is still required before the intent can be performed.
4 Analysis
4.1 Diversity of Slots and Intents
The generated LUCID data contains more intents and slots than existing task-oriented dialogue datasets (Table 1). Specifically, the dataset contains 100 intents, across 13 domains, with 501 different slots. While the SGD dataset contains more domains than LUCID, these domains are narrower in scope. For example, SGD includes separate domains for buses, taxis, flights and trains, while LUCID has a single transportation domain incorporating intents for each of these areas. The additional number of slots and intents in LUCID illustrates our ability to create diverse and challenging data using LUCID, despite generating a smaller dataset compared to SGD and MultiWOZ (see Table 4). Considering the size of our dataset, compared to SGD and MultiWoz there are almost twice as many unique slot values, demonstrating the level of variety in our dataset (see values per turn in Table 4).
We compare the number of slots within LUCID to PRESTO, SGD and MultiWoZ in Table 1, showing that after excluding duplicated slots in both PRESTO and SGD, the LUCID generated data has almost twice the number of slots. Appendix F outlines our approach for removing these duplicated slots, in the case of SGD when the same slots are included in different services, or in PRESTO when slots are counted multiple times if they have slightly different parse trees.
# Domains # Intents Ints per Dom # Slots # Labelled Unhappy Paths.
PRESTO - 34 - 303$\dagger$ 6
PRESTO-no dup - 34 - 276$\dagger$ 6
SGD 20 88 4.4 365 0
SGD-no dup 20 46 2.3 240 0
MultiWOZ 7 11 1.6 35 0
LUCID 13 100 7.7 501 9
Table 1: Summary statistics of our dataset, displaying the number of domains and intents present, the number of intents per domain, the number of slots present, and the number of explicitly labelled conversational phenomena (unhappy paths). $\dagger$ considered the 303 slots in the English intents. Unlike Table 4, this table considers all splits in the dataset.
It is worth pointing out that the LUCID dataset was generated primarily to showcase the capabilities of the LUCID data generation system. Others are free to use the LUCID system to generate much larger and even more complex datasets, with many more domains, intents and interesting conversational phenomena. This extensibility is what most clearly distinguishes LUCID from these other data generation efforts.
4.2 Conversational Phenomena
LUCID contains a greater number of labeled conversational phenomena than existing dialogue datasets (Table 1). The recently released PRESTO dataset also contains turn-level annotated phenomena, labelling six types of unhappy paths (Goel et al., 2023). These unhappy paths include in-turn corrections, correcting actions, correcting slot values, code-mixing, disfluencies and cancellations. While half of these phenomena relate to corrections, only two of our labeled phenomena are for corrections, correcting slot values either in-turn or across turns. Instead, we focus on distinguishing between relevant, sensible user replies and where a virtual assistant should ask for clarification rather than using the initial response to populate slots.
While not explicitly stated, some specific conversational phenomena can also be derived from existing datasets, such as co-references Cheng et al. (2020), or user corrections Rastogi et al. (2020).
4.3 Qualitative and Quantitative Analysis
We perform a qualitative analysis on the generated dataset, identifying issues that can be addressed by filtering out more examples in a post-processing step. Informed by this analysis, we perform a further quantitative analysis of 200 conversations in our dev set, measuring the quality of the system labels and investigating the impact of any identified issues. See Appendix A for the full results of our quantitative analysis, and Appendix B for details of additional post-processing. By publishing these more detailed quality metrics, we aim to raise the bar for future data generation efforts.
Finding 1) The natural language responses from the model do not always reflect the system labels that have been predicted. We observed that a correct system label can be accompanied by a natural language response that does not reflect the correct system prediction. We noticed this for sarcastic responses, where only the system label and not the natural language response reflected the user’s sarcasm. We choose to manually review the turns labelled as sarcastic, filtering out 4 dialogues. However, our quantitative analysis on the dev set highlights this as an issue beyond sarcastic turns, with natural language responses not faithfully following the system label predictions in 1% of conversations.
As a related issue, the natural language response can also suggest that an intent has been performed before confirmation is given by the user. Informed by this finding, we filter out conversations when an intent was not performed, unless there was a cancellation signal provided by the user (removing 79 conversations). After this filtering, our quantitative analysis finds that 4.5% of conversations include responses that suggest an intent has been performed before confirmation is given. However, in each case there was no impact on the conversation beyond the phrasing of the natural language response. As LUCID prioritises validating the system labels, there is room for improvement in future work to also improve the corresponding natural language responses.
Finding 2) The planner’s choice of slots and their corresponding values can sometimes be unrealistic. While a strength of LUCID is the realistic and varied slot values used in conversations, this is not always the case. We also notice that the choice of slots included in a conversation is not always realistic. For example, you would not usually give the start time, end time in addition to specifying the duration of a swimming lesson. Our quantitative analysis identifies that 6% of conversations contain at least one unrealistic slot value, while 3.5% of conversations include an unrealistic combination of slots. The unrealistic slot combinations demonstrate a limitation to our sampling approach, where we randomly sample which optional slots should be included in each conversation.
Finding 3) The user does not always understand the purpose of the intent. For example the user may ask ‘can you find my favorites from yesterday?’, when it is not clear if the user understands what a ‘favorite’ is. This is a consequence of our conversation plans telling users which intent should be performed, without also providing a description. The quantitative analysis finds that in 6.5% of conversations, either the user or the natural language response does not fully understand an intent, suggesting that descriptions should be included for future data generation work.
Finding 4) The system command labelling is consistently high quality, with few labelling mistakes. In a quantitative analysis of 200 conversations in the development set, we find only 1% of examples when the system label is not correct. In comparison, Eric et al. (2020) identify annotation errors in 40% of turns in MultiWoz 2.0, demonstrating the relative quality of the LUCID system labels. More detail on the two system labelling errors identified are provided in Appendix A.
Additionally, we find that only 8.5% of conversations have a string slot value that does not match the reviewer’s first choice. However, these differences are subjective and subtle, with over half of the cases concerning whether to include ‘the’ before a date (e.g. ‘the 5th of March’ vs ‘5th of March’).
5 Baseline Results
We train T5 Raffel et al. (2020) and Flan-T5 Chung et al. (2022) baseline models on LUCID, evaluating on both our in-distribution and out-of-distribution test sets. When retrieving intent schemas, a Sentence-BERT Reimers and Gurevych (2019) model is used to encode the tool name and the last user utterance, choosing the tool with the highest cosine similarity. As gold system labels for previous turns are seen in the prompt conversation history, we retrieve all tools that have been mentioned in an oracle history up to that point.
As expected, the joint goal accuracy is considerably higher when evaluating on the seen test set compared to the unseen test set, with accuracy scores of 67.5% and 42.2% respectively for a T5-base model (see Table 5). We also isolate the impact of the retrieval model, comparing three scenarios: 1) using our tool retrieval, 2) using an oracle tool retrieval, and 3) including no tools in the prompt. We find that the tool retrieval is not a major weakness of our baseline model (see Table 6).
Finally, we evaluate our Flan-T5-base model on each different conversational phenomena (see Section 3.5), highlighting sarcasm, ASR-early end examples, and answering about a different slot as the most challenging phenomena. On the other hand, cancellations, users delaying giving confirmation, and overheard answers are more easily predicted, and our baseline model makes no mistakes in these cases. Full details of our evaluation metrics, training setup and prompts can be found in Appendix C.
6 Conclusion
We introduce LUCID, a pipeline of LLM calls which is designed to create high quality and linguistically sophisticated dialogue data. LUCID involves an extensive validation process, including three validator LLMs that discard conversations where there is any disagreement. To demonstrate the quality of the data produced, we generate a seed dataset of 4,277 dialogues, consisting of 92,699 turns, with a wide variety of challenging conversational phenomena. These phenomena are all labeled at a turn level. The generated system labels in LUCID prove to be highly accurate, with only 1% of conversations containing system errors. This allows for reliable evaluation across our nine different conversational phenomena, either on seen or unseen intents. We also provide a detailed analysis of our generated data, identifying areas that could be improved in future data generation efforts. Finally, we make our code available to facilitate larger scale, high quality data generation with more intents, domains and complex conversational phenomena.
Appendix A Quantitative Analysis System Label Noise
Table 2 provides full results from our quantitative analysis of 200 dev examples. In this section, we additionally explain the issues that caused the two incorrect system labels identified.
In the first instance, the user says “I want to review the film the godfather. I give it a 9 out of 10 and my review is an absolute classic! Great performances and storytelling.. No spoilers in my review.” The system interprets this as giving the film name (the godfather), the rating (9), the spoiler alert (False), and the review text. However, the system also predicts the text ‘No spoilers in my review’ as part of the review text, when this may not be the case.
In the second instance, the user performs an in-turn correction, changing the value of the additional notes slot for a new hair appointment intent. However, the system predicts this correction as part of the note itself, giving the slot value: ‘I need a quick haircut, actually make that I’m getting ready for a family reunion photoshoot and want a new haircut.’
Appendix B Additional Post Processing
We provide more detail on the other filters that were used in post processing to improve the quality of the data.
We filter conversations where a slot being is corrected during the conversation, but there is no correctional signal provided by the user, used for our complex phenomena validation (mentioned in Section 3.5). This filters out instances where a slot was first mentioned by the user without giving a value, with the system incorrectly assigning a slot value from this turn. The user will then follow-up in another turn with the desired slot value. This new validation filters out 123 conversations, which entirely resolves this issue.
We perform additional filtering to remove empty string slot values (removing 27 conversations), and any instances where the system turn predicts a hint, as hints should only occur in Signal turns. There were 172 occurrences when a hint was predicted by the system, although in almost all cases these conversations were already filtered by another post-processing filter.
Issue name Prevalence
NLG indicates intent performed before confirmation 4.5%
NLG does not follow system prediction 1%
Unrealistic slot values 6%
Unrealistic combination of slots mentioned for an intent 3.5%
First choice string span not selected 8.5%
User or NLG does not fully understand purpose of the intent 6.5%
Table 2: Quantitative analysis from 200 dialogues in our development set. We report these six metrics in addition to the system label accuracy figure of 1% provided in Section 4.3.
Appendix C Baseline results
We train six different baseline models on the LUCID training data (T5-small, T5-base, T5-large, Flan-T5-small, Flan-T5-base and Flan-T5-large models). Each model is evaluated on the test set for seen intents and the OOD test set for unseen intents (see Table 5). We additionally experiment with training our Flan-T5-base baseline on varying amounts of training data (see Table 7). Details about about the choice of hyper-parameters can be found in Appendix I.
To measure performance on our generated LUCID data, we consider five performance metrics: joint goal accuracy, intent accuracy, fuzzy slot accuracy, exact match accuracy between user turns, and exact match accuracy for an entire dialogue.
For joint goal accuracy, we consider a fuzzy matching score for string slot values. As many system turns involve a say command, a joint goal accuracy figure is only calculated for turns where a value is predicted (or contained in the system labels). Additionally, we consider the goal state of all intents in the conversation, rather than considering different states for different intents or domains separately.
We also use fuzzy matching for our slot accuracy measure, which is a joint accuracy measure across all the slot values provided in a single system turn (when any slot values are predicted, or when they are included in the labels). We additionally introduce exact match metrics that consider the accuracy of all system commands, not just those that refer to intent and slot values (for example, including ‘say’ commands). We introduce two exact match metrics - exact match (turn) considers whether all predicted system commands between two user turns exactly match their labels, while exact match (conversation) considers whether every predicted system command in a conversation matches with the system labels. We introduce the exact match metrics as a result of the consistently high quality string slot values in the generated data (see Section 4.3).
The exact match between user turns is measured for a Flan-T5-base model for each conversational phenomena, both in the seen and unseen (OOD) test sets (see Table 3). We use this metric because some conversational phenomena involve an assignment, which is then followed by a ‘say’ command. As predicting a ‘say’ command following an assignment is not challenging for the model, we find that using the exact match between user turns metric provides the fairest comparison. The most challenging phenomena for our Flan-T5-base model are ASR-early end, sarcasm and answering about another slot phenomena (see Table 3), although predictions for ASR-early end are substantially worse for the seen intents. A number of phenomena appear to be less challenging than examples with no unhappy paths (see ‘None’ in Table 3), particularly for the OOD test set. This occurs because many phenomena do not involve any slot assignment, which becomes more challenging in the OOD test set.
Our baseline models are fine-tuned using the following prompt: “You are a smart AI assistant who is responsible for writing system commands to describe what the user has asked for. Your job is to write the next system command based on the latest user turn, considering the conversation so far.” When using tool retrieval, the following text is added “Information about the following tools may help:”, before providing the retrieved intent alongside intents from the conversation history. Finally, a single in-context example is provided to the model (see our code for further details).
Conv. phenomena Total Train Dev Test Test-OOD Test Test-OOD
# # # # # Acc. Acc.
Cancellation 12 5 2 3 2 100 100
ASR-early end 58 41 7 7 3 43 100
Sarcasm 63 46 3 8 6 75 67
Delay confirmation 76 53 7 4 12 100 100
Answer about another slot 113 75 13 11 14 64 43
Irrelevant answer 163 116 18 15 14 93 93
Overheard answer 203 153 17 23 10 100 100
In-turn correction 215 145 27 25 18 80 72
Correction 250 166 28 31 25 90 81
None 3,200 2,279 307 252 362 82 56
Conv. w/ 1+ unhappy path 1,077 754 108 119 96 - -
Total conversations 4,277 3,033 415 371 458 - -
% Conversations unhappy 25.2% 24.9% 26.0% 32.1% 21.0% - -
Table 3: Total number of each conversational phenomenon within each split of our dataset. While there are few conversations for cancellation, this behaves similarly to the ‘delay confirmation’ phenomenon. We also show the exact match (Turn) metric for each conversational phenomena from a T5-Flan-base baseline model (details of the metric are provided in section Appendix C.
Appendix D Intent generation
In total, 54 intent descriptions were provided, with a single intent removed for data quality reasons. The removed intent involved the user asking the virtual assistant to start watching a television channel, giving specific start and end times for when they want to start watching the channel. As this is an unrealistic scenario (a user would want to start watching a television channel straight away), the intent was removed.
Additionally, query intents were merged together for intents that had the same generated entity names. As a result, there are 6 fewer query intents than transactional intents, resulting in a total of 100 intents. Each of the following pairs of transactional intents returned the same entity names, and so their corresponding query intents were combined: add_tv_program_to_favorites and add_artist_to_favorites (both of which return ‘favorites’ entities), set_timer and set_alarm (both of which return ‘alarm’ entities), book_nails_appointment and book_spa_appointment (both of which return ‘appointments’), order_supermarket_shop and order_takeaway (both of which return ‘orders’), add_song_to_favorites and play_song (both of which return ‘songs’), and review_film and review_restaurant (both of which return ‘reviews’).
Appendix E Slot Value Generation
This section provides more detail to describe how the conversation planner creates slot values, using a range of different stages.
The slot values for the first intent are initially sampled based on the plausible slot values generated for each intent (see stage 2), preventing LLMs generating repetitive conversations. An LLM is then asked to make the slots for each intent more realistic and coherent (stage 4). This prevents contradictory slot values, for example when a hotel check-out date is before the check-in date. This stage also prevents highly unlikely slot values being over represented, while introducing further variety into the slot values being provided.
For every additional intent in the conversation, an LLM generates a plausible reason why the user would want to complete this intent given what has already occurred in the conversation (stage 5). Likely slot values are then generated (stage 6) based on this context. Stage 4 is then repeated, increasingly the likelihood that the slot values selected are realistic and coherent for the intent. An LLM is then asked to update the slot values across every intent in the conversation so that these intents are related and consistent (stage 7), encouraging more natural conversations that align more closely with human to virtual assistant interactions.
Each of the 100 intents used for our data generation are listed in Table 8 and Table 9. These tables list each transactional intent, along with its corresponding query intent. We also provide the human authored descriptions for each intent that were initially provided to LUCID.
Appendix F Slot duplication within PRESTO and SGD
SGD report 214 slots in their training data Rastogi et al. (2020), corresponding to 365 slots across all dataset splits (see Table 1). This counts slots with exactly the same names in the same domains within different services, which we consider to be duplicated slots (although the allowed slot values may change in each case). As a result, we provide a more direct comparison to SGD without this slot and intent duplication across services (see ‘SGD-no dup’ in Table 1).
For PRESTO, we consider the 303 slots present in the English split of the dataset (Goel et al., 2023). However, as the semantic annotations in PRESTO are represented in parse-trees, slots are counted multiple times if their paths are different. We find the number of English slots in PRESTO reduces to 276 without this duplication.
When considering the total number of slots in PRESTO, the same slot can be counted multiple times depending on its position in the labelled parse trees. For example, the Send_digital_object intent includes bcc and cc slots. Both of these slots can be a Personal_contact entity, which contains a person slot. In this case, the person slot within Personal_contact would be counted twice if it was contained within either the bcc or cc slots. Removing this slot duplication reduces the number of English slots in PRESTO from 303 to 276.
Note, we consider the v.2.2 of MultiWOZ for our comparison, as this version explicitly states the intents present in the dataset.
Appendix G Example Dialogues
In addition to the examples provided in Figure 1, Figure 3 and Figure 4, we provide three additional examples of the LUCID generated conversations. To provide an unbiased sample of our conversations, we show the first three dialogues in the dataset (see Figure 5, Figure 6 and Figure 7).
Appendix H Breakdown of issues identified by the LLM validation stages
In total, the two LLM validation agents described in stage 13 identified issues for 2,457 and 2,477 of turns, compared to the 557 issues identified by the LLM validation agent in stage 14 with access to the conversation rules. While the agent with access to these rules identifies fewer errors, a larger proportion of the issues identified are not identified by another validation agent (23%, compared to 7% for each other validation agent).
Appendix I Modelling setup, parameters, computing setup
For each baseline experiment, we train for 3 epochs. This was selected as a hyper-parameter based on the development set loss after training for 1-10 epochs for our Flan-T5-base baseline. We use a learning rate of $\times 10^{-5}$, with a linear learning schedule, a batch size of 2 with 8 gradient accumulation steps. For experiments with reduced training data, we train for more epochs as the training data is increased (with epochs inversely proportional to the size of the training data, allowing for a fair comparison).
Our baseline models have the following number of parameters: T5-small and Flan-T5-small (60 million parameters), T5-base and Flan-T5-base (220 million parameters), T5-large and Flan-T5-large (770 million parameters). We train our models with V100 GPUs, with our combined baseline experiments training for approximately 80 GPU hours.
Figure 6: An example conversation from LUCID (Example #2). As described in Appendix G, we show the first three examples to provide an unbiased sample from our generated data.
Figure 7: An example conversation from LUCID (Example #3). As described in Appendix G, we show the first three examples to provide an unbiased sample from our generated data.
DSTC2 WOZ2.0 FRAMES M2M MultiWOZ SGD LUCID
# domains 1 1 3 2 7 16 12
# dialogues 1,612 600 1,369 1,500 8,438 16,142 3,033
# turns 23,254 4,472 19,986 14,796 113,556 329,964 65,217
Turns per dialogue 14.49 7.45 14.60 9.86 13.46 20.44 21.50
No. of slots 8 4 61 13 24 214 432
No. of slot values 212 99 3,871 138 4,510 14,139 4,701
Values per turn 0.009 0.02 0.2 0.009 0.04 0.04 0.07
Table 4: Reported statistics for LUCID, and related datasets for task-oriented dialogue. All statistics refer to the training split of the datasets, except for Frames which reports figures for all splits. Compared to previous dialogue datasets, LUCID has considerably more slots, and more turns per dialogue. There are also more possible slot values per turn than either MultiWoZ or SGD.
Intent acc. Joint goal acc. Slot acc. Match (turn) Match (conv.)
Test (seen):
T5-Small 94.7 57.1 69.8 74.5 30.5
T5-Base 97.9 67.5 76.6 82.1 44.2
Flan-T5-Base 97.9 69.7 77.6 82.6 45.8
T5-Large 98.7 69.0 77.9 83.2 46.9
Flan-T5-Large 98.5 69.7 78.5 83.5 47.4
Test-OOD (unseen):
T5-Small 95.3 22.0 38.0 46.2 6.3
T5-Base 97.6 42.2 61.4 58.5 10.3
Flan-T5-Base 97.6 41.3 61.2 57.0 7.6
T5-Large 98.8 45.7 64.1 60.2 11.4
Flan-T5-Large 98.6 53.2 66.6 59.9 10.3
Table 5: Results of our baseline model trained for 3 epochs, using a SentenceBERT retrieval model. Each evaluation metric is described in more detail in Appendix C. Results are from a single seed in each case.
Intent acc. Joint goal acc. Slot acc. Match (turn) Match (conv.)
Test (seen):
No tools 97.4 66.1 75.9 81.6 43.7
w/ retrieval 97.9 69.7 77.6 82.6 45.8
Oracle tools 99.1 69.3 78.1 83.1 45.8
Test-OOD (unseen):
No tools 87.1 32.8 55.7 53.2 8.3
w/ retrieval 97.6 41.3 61.2 57.0 7.6
Oracle tools 99.4 40.8 61.3 57.4 6.6
Table 6: Results of a T5-Flan-base model with our tool retrieval, using oracle tools, and with no tools provided in the prompt. Each evaluation metric is described in more detail in Appendix C. Results are from a single seed in each case.
# Training ex. Intent acc. Joint goal acc. Slot acc. Match (turn) Match (conv.)
Test (seen):
125 88.8 29.3 49.1 57.1 10.2
250 91.0 37.5 57.7 65.2 15.6
500 91.9 51.2 64.9 70.6 22.6
1k 94.3 58.9 70.5 75.2 29.1
2k 96.4 62.6 73.8 78.5 37.2
4k 96.7 65.0 75.0 80.3 40.4
8k 97.3 66.8 76.2 81.1 42.3
16k 97.6 66.4 76.5 81.9 43.9
Full (24,786) 97.9 69.7 77.6 82.6 45.8
Test-OOD (unseen):
125 92.7 16.7 32.3 34.9 4.1
250 93.5 20.5 43.9 44.6 4.8
500 95.5 26.3 49.7 48.5 6.1
1k 96.7 34.3 56.4 54.2 9.6
2k 96.9 32.0 56.5 53.9 5.9
4k 97.0 37.0 59.5 57.0 6.8
8k 97.2 35.5 58.6 56.5 8.5
16k 97.4 32.7 57.2 55.9 7.0
Full (24,786) 97.6 41.3 61.2 57.0 7.6
Table 7: Results of a T5-Flan-base model trained with varying amounts of training data (count of the system turns provided). Each evaluation metric is described in more detail in Appendix C.
Transactional intents (1-30) Corresponding query Intent description
add_artist_to_favorites find_favorites Add artist to favourites
add_event find_events Add event
add_payment_card find_payment_cards Add payment card
add_restaurant_to_favorites find_favorite_restaurants Add restaurant to favourites
add_song_to_favorites find_songs Add song to favourites
add_to_favourites find_favourite_pages Add a page to favourites
add_tv_program_to_favorites find_favorites Add a TV program to favourites
add_user find_users Add user with access to calendar
block_sender find_blocked_senders Block sender
book_bus_ticket find_bus_tickets Book a bus ticket
book_city_tour find_city_tours Book a city tour
book_cruise find_cruises Book cruise
book_flight find_flights Book a flight
book_guide find_guides Book a guide
book_hair_appointment find_hair_appointments Book hair appointment
book_hotel_room find_hotel_rooms Book a hotel room
book_massage find_massages Book a massage
book_nails_appointment find_appointments Book appointment to do nails
book_pedicure find_pedicures Book a pedicure
book_spa_appointment find_appointments Book a spa appointment
book_swimming_lesson find_lessons Book swimming lesson
book_taxi find_taxis Book a taxi
book_train_journey find_train_journeys Book a train journey
book_triathlon find_triathlons Book triathlon
buy_film_tickets find_film_tickets Buy film tickets
create_direct_debit find_direct_debits Create direct debit
create_playlist find_playlists Create playlist
create_reminder find_reminders Create a reminder
create_workout_regime find_workouts Create workout regime
log_exercise find_exercises Log exercise
Table 8: Each transactional intent (1-30), alongside its respective query intent, and the description provided to LUCID that was used to generate the intent.
Transactional intents (31+) Corresponding query Intent description
make_song_recommendation find_recommendations Make song recommendation
open_web_page find_web_pages Open an internet page in a web browser
order_coffee find_coffee_orders Order coffee
order_supermarket_shop find_orders Order supermarket shop
order_takeaway find_orders Order takeaway
pay_bill find_bills Pay bill
play_audiobook find_audiobooks Play audiobook
play_film find_films Play film on streaming service
play_podcast_episode find_podcast_episodes Play a podcast episode
play_song find_songs Play a song
rent_accommodation find_accommodations Rent accommodation
rent_car find_cars Rent a car
reserve_table find_reservations Reserve a table
review_film find_reviews Review film
review_restaurant find_reviews Review a restaurant
send_email find_emails Send an email
send_invoice find_invoices Send invoice
send_message find_messages Send a message
set_alarm find_alarms Set an alarm
set_timer find_alarms Set a timer
set_volume find_volume Set the volume
transfer_money find_transactions Transfer money
write_note find_notes Write a note
Table 9: Each transactional intent (31+), alongside its respective query intent, and the description provided to LUCID that was used to generate the intent.