### v2

# LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues

Abstract
Spurred by recent advances in Large Language Models (LLMs), virtual assistants are poised to take a leap forward in terms of their dialogue capabilities. Yet a major bottleneck to achieving genuinely transformative task-oriented dialogue capabilities remains the scarcity of high quality data. Existing datasets, while impressive in scale, have limited domain coverage and contain few genuinely challenging conversational phenomena; those which are present are typically unlabelled, making it difficult to assess the strengths and weaknesses of models without time-consuming and costly human evaluation. Moreover, creating high quality dialogue data has until now required considerable human input, limiting both the scale of these datasets and the ability to rapidly bootstrap data for a new target domain. We aim to overcome these issues with LUCID, a modularised and highly automated LLM-driven data generation system that produces realistic, diverse and challenging dialogues. We use LUCID to generate a seed dataset of 4,277 conversations across 100 intents to demonstrate its capabilities, with a human review finding consistently high quality labels in the generated data [^1] Please see http://github.com/apple/ml-lucid-datagen for the data and project code.
LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues
Joe Stacey1thanks: Work undertaken while author was an intern at Apple         Jianpeng Cheng2        John Torr2        Tristan Guigue2 Joris Driesen2        Alexandru Coca3∗        Mark Gaynor2        Anders Johannsen2 1Imperial College London      2Apple            3University of Cambridge j.stacey20@imperial.ac.uk, ac2123@cam.ac.uk {jianpeng.cheng,jtorr,tguigue}@apple.com {joris_driesen,mgaynor,ajohannsen}@apple.com
1 Introduction
Figure 1: An extract of a LUCID conversation containing a challenging phenomenon. In this case, the second user response is most likely to be from an overheard conversation rather than providing the desired slot value.
As AI virtual assistants become more sophisticated, there is an increasing need for dialogue datasets with more challenging conversational phenomena for both fine-tuning and evaluation. Existing datasets include multi-turn, multi-intent and multi-domain conversations Rastogi et al. (2020); Budzianowski et al. (2018), in addition to multi-lingual datasets Goel et al. (2023); FitzGerald et al. (2023); Hung et al. (2022); Li et al. (2021). However, in each case, the number of intents covered is relatively small. Moreover, the conversational phenomena included in these datasets are often limited in scope [^2] The PRESTO dataset (Goel et al., 2023) does explicitly label specific conversational phenomena, but to the best of our knowledge it is unique in this respect. Additionally, current machine-to-machine data collection methods still involve varying degrees of human involvement, with humans paraphrasing machine generated templates into natural language, and/or manually crafting plausible sequences of intents as dialogue outlines (Shah et al., 2018; Rastogi et al., 2020).
To overcome these issues, we introduce LUCID, LLM-generated Utterances for Complex and Interesting Dialogues. LUCID is composed of a pipeline of modularised LLM calls that create realistic, accurate and complex data, allowing the data generation process to scale to more intents, slots and challenging conversational phenomena. LUCID involves automated intent generation, with a mock back-end [^3] The mock back-end converts intents into Python classes, which are then instantiated as objects created for each intent. This mock back-end then interacts with LLM-based user and system agents, generating dialogues without the need for human annotation.
Ensuring data quality is a central challenge for a machine-to-machine generation process. We address this issue by breaking down the generation process into a pipeline of multiple, simpler LLM calls, thereby compartmentalising the data generation task into manageable steps that an LLM can consistently perform accurately. In addition, we use multiple LLM-based validators which discard conversations that might contain an issue. Our if in doubt, discard philosophy ensures a high quality standard for the data being created.
We release the data generation code to enable large scale data generation across different intents and domains, with the option of adding additional, complex conversational flows. We also provide training data, validation data, and two tests sets, a test set for seen intents, and an additional test set for unseen intents, allowing for convenient out-of-distribution evaluation.
2 Related Work
2.1 Task Oriented Dialogue Datasets
The most popular approach for creating dialogue datasets involves human-to-human interactions, with user annotators interacting with Wizard of Oz (WoZ) annotators Budzianowski et al. (2018); El Asri et al. (2017); Zhu et al. (2020); Eric et al. (2017); Wen et al. (2017). While using human annotators can create diverse, large scale datasets, this is done at a considerable cost, with expert annotators required for accurate dialogue annotations. User annotators follow a generated conversation plan (Budzianowski et al., 2018; El Asri et al., 2017; Zhu et al., 2020), guiding their interactions with the WoZ agent. We find that even in a purely machine-to-machine setup, generating conversation plans for each dialogue remains an effective way to ensure conversational variety.
2.2 Automated Data Collection Methods
To reduce the workload of annotators, dataset collection is becoming increasingly automated. A popular approach is to generate conversation outlines, which are then converted into natural language by annotators Shah et al. (2018); Rastogi et al. (2020); Lin et al. (2021) or using natural language templates (Bordes et al., 2017). As these conversation outlines are simulated based on hard-coded rules, this can limit the diversity of the user behaviour.
Human involvement in automated data generation includes ensuring the quality of the dataset, paraphrasing user and agent responses (Shah et al., 2018; Rastogi et al., 2020), providing semantic annotations (Goel et al., 2023; Budzianowski et al., 2018), outlining the sequences of user intents (Rastogi et al., 2020), and identifying out of scope or incoherent examples (Goel et al., 2023). We show that, with recent advances in language modelling OpenAI (2023); Ouyang et al. (2022), by reducing the data generation task into manageable steps, and using our if in doubt, discard validation methodology, it is now possible to achieve the same quality in an almost entirely machine-to-machine generation process. Parallel work by Liu et al. (2024) also introduces an automated method for generating task-oriented dialogue data. While we focus on the accurate labelling of challenging and diverse conversations, Liu et al. (2024) consider a variety of user personas with different styles of system responses.
See Appendix A for related work generating data with LLMs for other tasks.
3 Method
LUCID decomposes the data generation process into 14 individual LLM calls, described here as stages, creating manageable steps that LLMs can perform accurately. Alongside our if in doubt, discard validation, reducing the complexity of each LLM call helps to ensure the quality of our generated data. The data generation process consists of four main components (see Figure 2): the generation of intents (stages 1-2), a conversational planner (stages 3-8), turn-by-turn generation of conversations (stages 9-12), and our validation process (stages 13 and 14). The turn-by-turn data generation involves a User LLM agent interacting with a System LLM agent, which in turn interacts with a mock back-end created for each intent.
Figure 2: The stages in the LUCID data generation, generating intents (stages 1-2), planning conversations (stages 3-8), generating the conversations (stages 9-12) and validating the system predictions (stages 13-14).
3.1 Intent Generation
Schema for each intent are generated by an LLM, using a short human-authored natural language description of the intent (stage 1) [^4] Our code also allows intent schema to be created manually. Using these descriptions, LUCID calls an LLM to generate the intent and slot names, as well as the data type of each slot and whether it is mandatory or optional. In total, 100 intents are generated across 13 domains (see Appendix H for a detailed breakdown of how transactional and query intents are generated).
The next stage (stage 2), involves generating plausible values for each slot. We use these slot values as a starting point for our conversation planner, helping to encourage varied conversations.
3.2 Conversation Planner
The conversation planner provides instructions that guide a user LLM agent down certain types of conversational flows. The planner specifies: 1) the sequence of intents, 2) the slot values for each intent, and 3) any complex conversational phenomena that must be included, specifying when and how these phenomena should be incorporated. This creates a plan that the user must adhere to, reducing the complexity of the data generation task, while also ensuring variety in the generated conversations. This plan is communicated to the User LLM at each turn through a series of conversation rules.
The planner also decides the sequence of intents that will be included in a conversation (stage 3). Depending on the primary intent used to start the conversation, the planner then decides which intents are likely to follow this intent, with the aim of creating both varied and realistic conversations. See Appendix B for further details about the planner.
3.2.1 Generating Slot Values
The slot values chosen by the planner substantially impact the conversations, and as a result, we have multiple, separate stages for generating the slot values (stages 4, 5, 6 & 7). This process involves updating the slot values to make sure these are realistic and coherent (stage 4), generating a reason why the user wants to perform any subsequent intents (stage 5), and generating slot values for the subsequent intents based on this justification (stage 6). Finally, an LLM updates the slot values across every intent in the conversation to ensure they are consistent and realistic when considered collectively (stage 7).
We additionally use an LLM to generate realistic entities to be returned after any queries (stage 8).
3.3 Generating Conversations
Conversations are generated turn-by-turn with a User LLM interacting with a System LLM, which in turn interacts (via Pythonic function calls and variable assignments) with a mock back-end for each intent. A Response LLM then communicates natural language responses back to the user. The user behaviour is governed by the conversational rules created by our planner, shaping the outcome of the conversations.
Conversations start with an utterance from the User LLM (stage 9), which is then interpreted and labelled by the System LLM (stage 10). Initially, no string slot values are predicted. These values are predicted in an additional stage (stage 11), where an LLM is instructed, where possible, to choose the string values from spans of the user utterance (avoiding hallucinations). The predicted semantic labels then interact with a mock back-end for each intent. The mock back-end then informs the System LLM about any missing slots or whether confirmation is required for the intent. Finally, the Response LLM responds back to the user (stage 12), requesting any additional slots or asking the user for confirmation.
3.4 LLM-based Validation
We implement an LLM-based validation process to ensure reliable and consistent labelling in the conversations, using our principle of if in doubt, discard. First, based on the observation that the system LLM is more uncertain about incorrect predictions, we repeat the system predictions twice (using a temperature value of 0.7), and abort the conversation if the three predictions are not identical (stage 13). Additional validation is then performed by another LLM (stage 14) which also labels the user requests, except this time with access to the conversation rules that the user is following. These predictions must also exactly match the original System LLM predictions, otherwise the conversation is aborted. The validation in stages 13 and 14 is performed before the string slot values are predicted, avoiding conversations being aborted when these slot values have slightly different phrasing. Further validation is described in Appendix E.
3.5 Introducing Additional Conversational Phenomena
To make interesting, diverse and challenging conversations, we introduce a wide range of conversational phenomena which are labelled automatically at a turn level (see Figure 4). These phenomena include sarcastic or irrelevant replies, or cases where the user is overheard in another conversation. LUCID also contains examples where a user corrects themselves, either within a turn (in-turn correction) or in a later turn (correction). Alternatively, a user may cancel an intent (cancellation) or delay confirming the intent until a future turn (delay confirmation). Our conversational phenomena also include cases where the virtual assistant requests a value for one slot, but the user responds about a different slot (respond different slot). Finally we also include ASR early end errors (ASR-early end), where the LLM produces truncated slot values where the user text is abruptly cut off. See Table 4 for the full distribution of these phenomena.
Figure 3: A (simplified) example labelled conversation. Each dialogue contains user, system, signal and response turns.
3.6 Annotation Scheme and our Mock Back-end
We apply a concise labelling system to track the states for each intent in a conversation. This labelling system follows a Pythonic syntax, with function calls used to initialise intents and entities when these are first mentioned, and attribute assignments used for any subsequent slot filling operations (see Figure 3). This schema-based, concise form of semantic labelling is highly convenient, and avoids the need for state tracking for each individual turn.
Our labelling involves four types of turns: 1) User turns; 2) System turns, labelling user intentions; 3) Signal turns, returned after our mock back-end processes the system command; and 4) Response turns, which are natural language responses back to the user. Further details on our labelling schema is provided in Appendix F.
4 Analysis
4.1 Diversity of Slots and Intents
The generated LUCID data contains more intents and slots than existing task-oriented dialogue datasets (Table 1). Specifically, the dataset contains 100 intents, across 13 domains, with 501 different slots. While the SGD dataset contains more domains than LUCID, these domains are narrower in scope. For example, SGD includes separate domains for buses, taxis, flights and trains, while LUCID has a single transportation domain incorporating intents for each of these areas. The larger number of slots and intents in LUCID illustrates our ability to create diverse and challenging data using LUCID, despite generating a smaller dataset compared to SGD and MultiWOZ (see Table 5).
# Domains # Intents Ints per Dom # Slots # Labelled Unhappy Paths.
PRESTO - 34 - 303$\dagger$ 6
PRESTO-no dup - 34 - 276$\dagger$ 6
SGD 20 88 4.4 365 0
SGD-no dup 20 46 2.3 240 0
MultiWOZ 7 11 1.6 35 0
LUCID 13 100 7.7 501 9
Table 1: Summary statistics of our dataset, displaying the number of domains and intents present, the number of intents per domain, the number of slots present, and the number of explicitly labelled conversational phenomena (unhappy paths). For PRESTO, we consider the 303 slots in English intents ($\dagger$). Unlike Table 5, this table considers all splits in the dataset. Appendix I describes how duplicate slots and intents are removed for SGD and PRESTO.
As the LUCID dataset was generated primarily to showcase the capabilities of the LUCID data generation system, others are free to use the LUCID system to generate much larger and even more complex datasets. This extensibility is what most clearly distinguishes LUCID from these other data generation efforts.
4.2 Conversational Phenomena
LUCID contains a greater number of labelled conversational phenomena than existing dialogue datasets (Table 1). The recently released PRESTO dataset also contains turn-level annotated phenomena, labelling six types of unhappy paths (Goel et al., 2023). These unhappy paths include in-turn corrections, correcting actions, correcting slot values, code-mixing, disfluencies and cancellations. While half of these phenomena relate to corrections, this is the case for only two of our labelled phenomena, correcting slot values either in-turn or across turns. Instead, we focus on distinguishing between relevant, sensible user replies from cases where a virtual assistant should ask for clarification (rather than using the initial response to populate slot values).
4.3 Qualitative and Quantitative Analysis
We perform a qualitative analysis on the generated dataset (conducted by one of the paper authors) to thoroughly investigate the dataset quality and identify any issues. This included a manual review of 200 conversations in our dev set, which only highlighted two labelling errors (impacting only 1% of conversations). In comparison, Eric et al. (2020) identify annotation errors in 40% of turns in MultiWoz 2.0, demonstrating the relative quality of the LUCID system labels.
The two labelling errors identified in this review involved: 1) A user mentioning there will be no spoilers in a review, where LUCID correctly assigns the spoiler alert slot value as False, but additionally includes the text ‘no spoilers in my review’ as part of the review itself. 2) LUCID not recognising an in-turn correction by the user, mistakenly including all of the user text (including the correction itself) as part of a slot value. See Appendix C for further details. We additionally share a qualitative analysis of our data in Appendix D, highlighting specific areas where our method could be further improved. We include this analysis to further raise the bar for future LLM data generation efforts.
5 Baseline Results
Intent acc. JGA
Test (seen):
T5-Small 94.7 57.1
T5-Base 97.9 67.5
T5-Large 98.7 69.0
Test-OOD (unseen):
T5-Small 95.3 22.0
T5-Base 97.6 42.2
T5-Large 98.8 45.7
Table 2: Results of our baseline model (with retrieval). Full results and evaluation metric descriptions are provided in Appendix G.
We train T5 Raffel et al. (2020) and Flan-T5 Chung et al. (2022) baseline models on LUCID, evaluating on both our in-distribution and out-of-distribution test sets. When retrieving intent schemas, a Sentence-BERT Reimers and Gurevych (2019) model is used to encode the tool name and the last user utterance, choosing the tool with the highest cosine similarity.
As expected, the joint goal accuracy is considerably higher when evaluating on the seen test set compared to the unseen test set, with accuracy scores of 67.5% and 42.2% respectively for a T5-base model (see Table 6). We also isolate the impact of the retrieval model, comparing three scenarios: 1) using our tool retrieval, 2) using an oracle tool retrieval, and 3) including no tools in the prompt. We find that the tool retrieval is not a major weakness of our baseline model (see Table 7). Finally, we evaluate our Flan-T5-base model on each different conversational phenomena (see Section 3.5), highlighting sarcasm, ASR-early end examples, and answering about a different slot as the most challenging phenomena. Full experimentation details and results can be found in Appendix G.
6 Conclusion
We introduce LUCID, a pipeline of LLM calls which is designed to create high quality and linguistically sophisticated dialogue data. LUCID involves an extensive validation process, including three validator LLMs that discard conversations where there is any disagreement. To demonstrate the quality of the data produced, we generate a seed dataset of 4,277 dialogues, consisting of 92,699 turns, with a wide variety of challenging conversational phenomena. The generated system labels in LUCID prove to be highly accurate, with only 1% of conversations containing a labelling error. We make our code available to facilitate larger scale, high quality data generation.
Limitations
The main limitation of our approach is the cost of using a closed-source LLM. This prevented us from generating a larger number of dialogues or performing more ablation studies to isolate the improvements from specific stages. This cost was driven by our if in doubt, discard approach to validation, which prioritised the accuracy and quality of the data produced, at the expense of the computational time and cost involved. While there are also substantial costs associated with high quality manual annotation, in this work we aim to show that an LLM-driven approach to generating high quality data is possible and feasible. We also aim to produce a seed dataset of the highest quality which can be used by practitioners on an on-going basis.
Appendix A Related work - Data Generation with LLMs
Wu et al. (2023) recently introduce a framework allowing the interaction of multiple, different LLMs, based on the idea that LLMs can solve highly challenging tasks if these tasks are broken into smaller steps. While Wu et al. (2023) are successful in generating dialogues for a group chat scenario, this does not require the intent and slot labelling needed for task-oriented dialogue. For other NLP tasks, to avoid labelling errors or poor quality data, generating data with LLMs can involve human annotators reviewing the generated utterances Liu et al. (2022); Wiegreffe et al. (2022), or using the generated data as unlabelled data to be used with knowledge distillation Meng et al. (2022); He et al. (2022); Stacey and Rei (2023). Labelled data generation has been successful for other tasks without human input Honovich et al. (2023); Ye et al. (2022); Schick and Schütze (2021); West et al. (2022); Wu et al. (2022), however noise may be an issue for a large proportion of the data Honovich et al. (2023); Schick and Schütze (2021).
Appendix B Conversation Planner Details
To ensure variety in the generated conversations, the planner makes extensive use of sampling, choosing how many intents should be provided in the conversation, which optional slots should be discussed, the conversational phenomena (both happy and unhappy paths) that should be included, and which slots and intents any conversational phenomena should be applied to.
Sampling is also involved in choosing the path of intents that are included in the conversation. The planner is provided with a sample of intents across all domains, before an LLM generates a plausible sequence of intents from this sample (stage 3).
We also provide more detail below to describe exactly how the conversation planner creates slot values for the conversation, involving a range of different stages:
The slot values for the first intent are initially sampled based on the plausible slot values generated for each intent (see stage 2), preventing LLMs generating repetitive conversations. An LLM is then asked to make the slots for each intent more realistic and coherent (stage 4). This prevents contradictory slot values, for example when a hotel check-out date is before the check-in date. This stage also prevents highly unlikely slot values being over represented, while introducing further variety into the slot values being provided.
For all other intents in the conversation (after the first intent), an LLM generates a plausible reason why the user would want to complete this intent given what has already occurred in the conversation (stage 5). Likely slot values are then generated (stage 6) based on this context. Stage 4 is then repeated, increasingly the likelihood that the slot values selected are realistic and coherent for the intent. An LLM is then asked to update the slot values across every intent in the conversation so that these intents are related and consistent (stage 7), encouraging more natural conversations that align more closely with human to virtual assistant interactions.
Appendix C Quantitative Analysis Findings
Table 3 provides the full results from our quantitative analysis of 200 dev examples. This analysis identified two labelling issues within the 200 conversations that were reviewed (conducted by one of the paper authors).
In the first instance, the user says “I want to review the film the godfather. I give it a 9 out of 10 and my review is an absolute classic! Great performances and storytelling.. No spoilers in my review.” The system interprets this as giving the film name (the godfather), the rating (9), the spoiler alert (False), and the review text. However, the system also predicts the text ‘No spoilers in my review’ as part of the review text, when this may not be the case.
In the second instance, the user performs an in-turn correction, changing the value of the additional notes slot for a new hair appointment intent. However, the system predicts this correction as part of the note itself, giving the slot value: ‘I need a quick haircut, actually make that I’m getting ready for a family reunion photoshoot and want a new haircut.’
Table 3 also measures some of the issues reported in the qualitative analysis, including: 1) how many times the natural language generated response (NLG) indicates an intent was performed before confirmation was given, 2) when the NLG does not follow the system predictions, 3) unrealistic slot values, 4) unrealistic combination of slots mentioned for an intent, 5) when the first choice string span was not selected by LUCID, and 6) when the user or NLG does not fully understand the purpose of the intent. While previous work does not report similar metrics, we publish these figures with the aim of raising the bar for future data generation efforts.
In the case of point 5) considering when the first choice string span was not selected, we find that only 8.5% of conversations have a string slot value that does not match the reviewer’s first choice. However, these differences are subjective and subtle, with over half of the cases concerning whether to include ‘the’ before a date (e.g. ‘the 5th of March’ vs ‘5th of March’).
Issue name Prevalence
NLG indicates intent performed before confirmation 4.5%
NLG does not follow system prediction 1%
Unrealistic slot values 6%
Unrealistic combination of slots mentioned for an intent 3.5%
First choice string span not selected 8.5%
User or NLG does not fully understand purpose of the intent 6.5%
Table 3: Quantitative analysis from 200 dialogues in our development set. We report these six metrics in addition to the system label accuracy figure of 1% provided in Section 4.3.
Appendix D Qualitative Analysis
We perform a qualitative analysis on our dev set, understanding potential limitations of our data, and suggesting ways these could be mitigated for future generation efforts.
Finding 1) The natural language responses from the model do not always reflect the system labels that have been predicted. We observed that a correct system label can be accompanied by a natural language response that does not reflect the correct system prediction. We noticed this for sarcastic responses, where only the system label and not the natural language response reflected the user’s sarcasm. We choose to manually review the turns labelled as sarcastic, filtering out 4 dialogues. However, our quantitative analysis on the dev set highlights this as an issue beyond sarcastic turns, with natural language responses not faithfully following the system label predictions in 1% of conversations.
As a related issue, the natural language response can also suggest that an intent has been performed before confirmation is given by the user. Informed by this finding, we filter out conversations when an intent was not performed, unless there was a cancellation signal provided by the user (removing 79 conversations). After this filtering, our quantitative analysis finds that 4.5% of conversations include responses that suggest an intent has been performed before confirmation is given. However, in each case there was no impact on the conversation beyond the phrasing of the natural language response. As LUCID prioritises validating the system labels, we do not implement validation checks on the natural language response. Introducing additional validation for the natural language responses is likely to also improve their quality.
Finding 2) The planner’s choice of slots and their corresponding values can sometimes be unrealistic. While a strength of LUCID is the realistic and varied slot values used in conversations, this is not always the case. We also notice that the choice of slots included in a conversation is not always realistic. For example, you would not usually give the start time, end time in addition to specifying the duration of a swimming lesson. Our quantitative analysis identifies that 6% of conversations contain at least one unrealistic slot value, while 3.5% of conversations include an unrealistic combination of slots. The unrealistic slot combinations demonstrate a limitation to our sampling approach, where we randomly sample which optional slots should be included in each conversation. This issue could be overcome with an additional LLM stage responsible for deciding if the slot combination provided is realistic or not.
Finding 3) The user does not always understand the purpose of the intent. For example the user may ask ‘can you find my favorites from yesterday?’, when it is not clear if the user understands what a ‘favorite’ is. This is a consequence of our conversation plans telling users which intent should be performed, without also providing a description. The quantitative analysis finds that in 6.5% of conversations, either the user or the natural language response does not fully understand an intent, suggesting that descriptions should be included for future data generation work.
Finding 4) The system command labelling is consistently high quality, with few labelling mistakes. We quantify this finding with our quantitative analysis of 200 conversations in the development set, which finds only 1% of examples when the system label is not correct. More detail on the two system labelling errors identified are provided in Appendix C.
Appendix E Validation and Post-Processing
We introduce additional validation, ensuring turns that include our challenging conversational phenomena are correctly predicted by our LLMs. When the conversation rules instruct a user to introduce a specific conversational phenomenon for a certain slot value, the user is instructed to also provide a signal (in the form of a special token) to show that this unhappy path is being performed. We then use this signal for validation purposes, ensuring that the following system command matches the expected response for this phenomenon. However, we do not provide these special tokens to the system which interprets and labels the user request; these would not be available to a real virtual assistant and we find that including them during data generation can result in unrealistic target labels (e.g. if a user’s ‘irrelevant’ response accidentally constitutes a plausible slot value).
A range of post-processing rules are also introduced after our qualitative analysis. We filter conversations where a slot is corrected during the conversation, but where there is no correctional signal provided by the user (as described above, a signal is provided by the user for each complex conversational phenomena which is used purely for validation purposes). This filtering process removes instances where a slot was first mentioned by the user without giving a value, with the system incorrectly assigning a slot value from this turn (filtering 123 conversations). We perform additional filtering to remove empty string slot values (removing 27 conversations), and any instances where the system turn predicts a hint, as hints should only occur in Signal turns. There were 172 occurrences when a hint was predicted by the system, although in almost all cases these conversations were already filtered by another post-processing filter.
In total, 56% of conversations pass all of our validation checks. To avoid wasting valuable conversational data, we salvage the prefix of an aborted conversation up to the point where the validation error was identified [^5] To avoid overly short conversations, we do this only if at least one intent has been performed already or at least 10 turns have occurred. In these cases, we truncate the conversation, sampling LLM generated natural language responses that justify interrupting the conversation.
Appendix F Annotation Schema Detail
An important part of our annotation schema is the order of the turns, and how system turns trigger the natural language responses. This section provides more detail on these points.
System turns always follow user turns. In most cases, the first system turn is followed by a signal turn, except when the system decides to immediately call a response with ‘say()’, for example if the user response is irrelevant. Signal turns are then followed by a system turn, which triggers the natural language response turn. The system turns therefore decide when to pass information to the mock back-end, and when to trigger the natural language response. We use the system turns as the targets in this dataset.
LUCID automatically creates the mock back-end for each intent using the schema generated in steps 1 and 2. This involves generating a Python class to represent the intent in question, which is then instantiated as an object and interacts with the system commands to indicate when mandatory slots have not yet been provided, or when confirmation is still required before the intent can be performed. The outputs of the mock back-end are represented by the signal turns described above.
Appendix G Baseline Results
We train six different baseline models on the LUCID training data (T5-small, T5-base, T5-large, Flan-T5-small, Flan-T5-base and Flan-T5-large models). Each model is evaluated on the test set for seen intents and the OOD test set for unseen intents (see Table 6). We additionally experiment with training our Flan-T5-base baseline on varying amounts of training data (see Table 8). Details about about the choice of hyper-parameters can be found in Appendix K.
To measure performance on our generated LUCID data, we consider five performance metrics: joint goal accuracy, intent accuracy, fuzzy slot accuracy, exact match accuracy (between user turns), and exact match accuracy for an entire dialogue.
For joint goal accuracy, we consider a fuzzy matching score for string slot values. As many system turns involve a say command, a joint goal accuracy figure is only calculated for turns where a value is predicted (or contained in the system labels). Additionally, we consider the goal state of all intents in the conversation, rather than considering different states for different intents or domains separately.
We also use fuzzy matching for our slot accuracy measure, which is a joint accuracy measure across all the slot values provided in a single system turn (when any slot values are predicted, or when they are included in the labels). We additionally introduce exact match metrics that consider the accuracy of all system commands, not just those that refer to intent and slot values (for example, including ‘say’ commands). We introduce two exact match metrics - exact match (turn) considers whether all predicted system commands between two user turns exactly match their labels, while exact match (conversation) considers whether every predicted system command in a conversation matches with the system labels.
The exact match between user turns is measured for a Flan-T5-base model for each conversational phenomena, both in the seen and unseen (OOD) test sets (see Table 4). We use this metric because some conversational phenomena involve an assignment, which is then followed by a ‘say’ command. As predicting a ‘say’ command following an assignment is not challenging for the model, we find that using the exact match between user turns metric provides the fairest comparison. The most challenging phenomena for our Flan-T5-base model are ASR-early end, sarcasm and answering about another slot phenomena (see Table 4), although predictions for ASR-early end are substantially worse for the seen intents. A number of phenomena appear to be less challenging than examples with no unhappy paths (see ‘None’ in Table 4), particularly for the OOD test set. This occurs because many phenomena do not involve any slot assignment, which becomes more challenging in the OOD test set.
For the tool retrieval, as gold system labels for previous turns are seen in the prompt conversation history, we retrieve all tools that have been mentioned in an oracle history up to that point.
Our baseline models are fine-tuned using the following prompt: “You are a smart AI assistant who is responsible for writing system commands to describe what the user has asked for. Your job is to write the next system command based on the latest user turn, considering the conversation so far.” When using tool retrieval, the following text is added “Information about the following tools may help:”, before providing the retrieved intent alongside intents from the conversation history. Finally, a single in-context example is provided to the model (see our code for further details).
Conv. phenomena Total Train Dev Test Test-OOD Test Test-OOD
# # # # # Acc. Acc.
Cancellation 12 5 2 3 2 100 100
ASR-early end 58 41 7 7 3 43 100
Sarcasm 63 46 3 8 6 75 67
Delay confirmation 76 53 7 4 12 100 100
Answer about another slot 113 75 13 11 14 64 43
Irrelevant answer 163 116 18 15 14 93 93
Overheard answer 203 153 17 23 10 100 100
In-turn correction 215 145 27 25 18 80 72
Correction 250 166 28 31 25 90 81
None 3,200 2,279 307 252 362 82 56
Conv. w/ 1+ unhappy path 1,077 754 108 119 96 - -
Total conversations 4,277 3,033 415 371 458 - -
% Conversations unhappy 25.2% 24.9% 26.0% 32.1% 21.0% - -
Table 4: Total number of each conversational phenomenon within each split of our dataset. While there are few conversations for cancellation, this behaves similarly to the ‘delay confirmation’ phenomenon. We also show the exact match (Turn) metric for each conversational phenomena from a T5-Flan-base baseline model (details of the metric are provided in section Appendix G).
Appendix H Intent Generation
In total, 54 intent descriptions were provided, with a single intent removed for data quality reasons. The removed intent involved the user asking the virtual assistant to start watching a television channel, giving specific start and end times for when they want to start watching the channel. As this is an unrealistic scenario (a user would want to start watching a television channel straight away), the intent was removed.
A transactional intent was created for each of the remaining 53 descriptions. LUCID then creates a query intent corresponding to each transactional intent. The query intent returns entities that would be created using the corresponding transactional intent.
Some query intents were merged together (this happens when the corresponding transactional intents had the same entity names - one of the intent properties generated by the LLM). As a result, there are 6 fewer query intents than transactional intents, resulting in a total of 100 intents. Each of the following pairs of transactional intents returned the same entity names, and so their corresponding query intents were combined: add_tv_program_to_favorites and add_artist_to_favorites (both of which return ‘favorites’ entities), set_timer and set_alarm (both of which return ‘alarm’ entities), book_nails_appointment and book_spa_appointment (both of which return ‘appointments’), order_supermarket_shop and order_takeaway (both of which return ‘orders’), add_song_to_favorites and play_song (both of which return ‘songs’), and review_film and review_restaurant (both of which return ‘reviews’).
Each of the 100 intents used for our data generation are listed in Table 9 and Table 10. These tables list each transactional intent, along with its corresponding query intent. We also provide the human authored descriptions for each intent that were initially provided to LUCID.
Appendix I Slot Duplication within PRESTO and SGD
SGD report 214 slots in their training data Rastogi et al. (2020), corresponding to 365 slots across all dataset splits (see Table 1). This counts slots with exactly the same names in the same domains within different services, which we consider to be duplicated slots (although the allowed slot values may change in each case). As a result, we provide a more direct comparison to SGD without this slot and intent duplication across services (see ‘SGD-no dup’ in Table 1).
For PRESTO, we consider the 303 slots present in the English split of the dataset (Goel et al., 2023). However, as the semantic annotations in PRESTO are represented in parse-trees, slots are counted multiple times if their paths are different. We find the number of English slots in PRESTO reduces to 276 without this duplication.
When considering the total number of slots in PRESTO, the same slot can be counted multiple times depending on its position in the labelled parse trees. For example, the Send_digital_object intent includes bcc and cc slots. Both of these slots can be a Personal_contact entity, which contains a person slot. In this case, the person slot within Personal_contact would be counted twice if it was contained within either the bcc or cc slots. Removing this slot duplication reduces the number of English slots in PRESTO from 303 to 276.
Note, we consider the v.2.2 of MultiWOZ for our comparison, as this version explicitly states the intents present in the dataset.
Appendix J Example Dialogues
In addition to the examples provided in Figure 1 and Figure 3, we provide three additional examples of the LUCID generated conversations. To provide an unbiased sample of our conversations, we show the first three dialogues in the dataset (see Figure 5, Figure 6 and Figure 7). We also show examples of each of the unhappy paths used in our dataset (see Figure 4).
Appendix K Modelling Setup, Parameters, Computing Setup
For each baseline experiment, we train for 3 epochs. This was selected as a hyper-parameter based on the development set loss after training for 1-10 epochs for our Flan-T5-base baseline. We use a learning rate of $\times 10^{-5}$, with a linear learning schedule, a batch size of 2 with 8 gradient accumulation steps. For experiments with reduced training data, we train for more epochs as the training data is increased (with epochs inversely proportional to the size of the training data, allowing for a fair comparison). For each LLM call, we use a temperature value of 0.7.
Our baseline models have the following number of parameters: T5-small and Flan-T5-small (60 million parameters), T5-base and Flan-T5-base (220 million parameters), T5-large and Flan-T5-large (770 million parameters). We train our models with V100 GPUs, with our combined baseline experiments training for approximately 80 GPU hours.
Figure 6: An example conversation from LUCID (Example #2). As described in Appendix J, we show the first three examples to provide an unbiased sample from our generated data.
Figure 7: An example conversation from LUCID (Example #3). As described in Appendix J, we show the first three examples to provide an unbiased sample from our generated data.
DSTC2 WOZ2.0 FRAMES M2M MultiWOZ SGD LUCID
# domains 1 1 3 2 7 16 12
# dialogues 1,612 600 1,369 1,500 8,438 16,142 3,033
# turns 23,254 4,472 19,986 14,796 113,556 329,964 65,217
Turns per dialogue 14.49 7.45 14.60 9.86 13.46 20.44 21.50
No. of slots 8 4 61 13 24 214 432
No. of slot values 212 99 3,871 138 4,510 14,139 4,701
Values per turn 0.009 0.02 0.2 0.009 0.04 0.04 0.07
Table 5: Reported statistics for LUCID, and related datasets for task-oriented dialogue. All statistics refer to the training split of the datasets, except for Frames which reports figures for all splits. Compared to previous dialogue datasets, LUCID has considerably more slots, and more turns per dialogue. There are also more possible slot values per turn than either MultiWoZ or SGD. The number of turns in LUCID refer to User, System, Signal and Response turns.
Intent acc. Joint goal acc. Slot acc. Match (turn) Match (conv.)
Test (seen):
T5-Small 94.7 57.1 69.8 74.5 30.5
T5-Base 97.9 67.5 76.6 82.1 44.2
Flan-T5-Base 97.9 69.7 77.6 82.6 45.8
T5-Large 98.7 69.0 77.9 83.2 46.9
Flan-T5-Large 98.5 69.7 78.5 83.5 47.4
Test-OOD (unseen):
T5-Small 95.3 22.0 38.0 46.2 6.3
T5-Base 97.6 42.2 61.4 58.5 10.3
Flan-T5-Base 97.6 41.3 61.2 57.0 7.6
T5-Large 98.8 45.7 64.1 60.2 11.4
Flan-T5-Large 98.6 53.2 66.6 59.9 10.3
Table 6: Results of our baseline model trained for 3 epochs, using a SentenceBERT retrieval model. Each evaluation metric is described in more detail in Appendix G. Results are from a single seed in each case.
Intent acc. Joint goal acc. Slot acc. Match (turn) Match (conv.)
Test (seen):
No tools 97.4 66.1 75.9 81.6 43.7
w/ retrieval 97.9 69.7 77.6 82.6 45.8
Oracle tools 99.1 69.3 78.1 83.1 45.8
Test-OOD (unseen):
No tools 87.1 32.8 55.7 53.2 8.3
w/ retrieval 97.6 41.3 61.2 57.0 7.6
Oracle tools 99.4 40.8 61.3 57.4 6.6
Table 7: Results of a T5-Flan-base model with our tool retrieval, using oracle tools, and with no tools provided in the prompt. Each evaluation metric is described in more detail in Appendix G. Results are from a single seed in each case.
# Training ex. Intent acc. Joint goal acc. Slot acc. Match (turn) Match (conv.)
Test (seen):
125 88.8 29.3 49.1 57.1 10.2
250 91.0 37.5 57.7 65.2 15.6
500 91.9 51.2 64.9 70.6 22.6
1k 94.3 58.9 70.5 75.2 29.1
2k 96.4 62.6 73.8 78.5 37.2
4k 96.7 65.0 75.0 80.3 40.4
8k 97.3 66.8 76.2 81.1 42.3
16k 97.6 66.4 76.5 81.9 43.9
Full (24,786) 97.9 69.7 77.6 82.6 45.8
Test-OOD (unseen):
125 92.7 16.7 32.3 34.9 4.1
250 93.5 20.5 43.9 44.6 4.8
500 95.5 26.3 49.7 48.5 6.1
1k 96.7 34.3 56.4 54.2 9.6
2k 96.9 32.0 56.5 53.9 5.9
4k 97.0 37.0 59.5 57.0 6.8
8k 97.2 35.5 58.6 56.5 8.5
16k 97.4 32.7 57.2 55.9 7.0
Full (24,786) 97.6 41.3 61.2 57.0 7.6
Table 8: Results of a T5-Flan-base model trained with varying amounts of training data (count of the system turns provided). Each evaluation metric is described in more detail in Appendix G.
Transactional intents (1-30) Corresponding query Intent description
add_artist_to_favorites find_favorites Add artist to favourites
add_event find_events Add event
add_payment_card find_payment_cards Add payment card
add_restaurant_to_favorites find_favorite_restaurants Add restaurant to favourites
add_song_to_favorites find_songs Add song to favourites
add_to_favourites find_favourite_pages Add a page to favourites
add_tv_program_to_favorites find_favorites Add a TV program to favourites
add_user find_users Add user with access to calendar
block_sender find_blocked_senders Block sender
book_bus_ticket find_bus_tickets Book a bus ticket
book_city_tour find_city_tours Book a city tour
book_cruise find_cruises Book cruise
book_flight find_flights Book a flight
book_guide find_guides Book a guide
book_hair_appointment find_hair_appointments Book hair appointment
book_hotel_room find_hotel_rooms Book a hotel room
book_massage find_massages Book a massage
book_nails_appointment find_appointments Book appointment to do nails
book_pedicure find_pedicures Book a pedicure
book_spa_appointment find_appointments Book a spa appointment
book_swimming_lesson find_lessons Book swimming lesson
book_taxi find_taxis Book a taxi
book_train_journey find_train_journeys Book a train journey
book_triathlon find_triathlons Book triathlon
buy_film_tickets find_film_tickets Buy film tickets
create_direct_debit find_direct_debits Create direct debit
create_playlist find_playlists Create playlist
create_reminder find_reminders Create a reminder
create_workout_regime find_workouts Create workout regime
log_exercise find_exercises Log exercise
Table 9: Each transactional intent (1-30), alongside its respective query intent, and the description provided to LUCID that was used to generate the intent.
Transactional intents (31+) Corresponding query Intent description
make_song_recommendation find_recommendations Make song recommendation
open_web_page find_web_pages Open an internet page in a web browser
order_coffee find_coffee_orders Order coffee
order_supermarket_shop find_orders Order supermarket shop
order_takeaway find_orders Order takeaway
pay_bill find_bills Pay bill
play_audiobook find_audiobooks Play audiobook
play_film find_films Play film on streaming service
play_podcast_episode find_podcast_episodes Play a podcast episode
play_song find_songs Play a song
rent_accommodation find_accommodations Rent accommodation
rent_car find_cars Rent a car
reserve_table find_reservations Reserve a table
review_film find_reviews Review film
review_restaurant find_reviews Review a restaurant
send_email find_emails Send an email
send_invoice find_invoices Send invoice
send_message find_messages Send a message
set_alarm find_alarms Set an alarm
set_timer find_alarms Set a timer
set_volume find_volume Set the volume
transfer_money find_transactions Transfer money
write_note find_notes Write a note
Table 10: Each transactional intent (31+), alongside its respective query intent, and the description provided to LUCID that was used to generate the intent.